The resistance of a platinum wire at 0◦C, 100◦C and the boiling point of

sulphur(444.6◦C) is found to be 2.56Ω, 3.56Ω and 6.78Ω respectively. The

resistance of the wire at a temperature t◦C is given by Rt=Ro(1+αt+βt2).

Calculate the ratio α/β.


1
Expert's answer
2021-05-23T18:55:11-0400
"R_0=2.56Ω\\3.56=2.56(1+\alpha(100)+\beta(100)^2)\\6.78=2.56(1+\alpha(444.6)+\beta(444.6)^2)"

"\frac{\alpha}{\beta}=-6880"